OpenStudy (gabbyalicorn):

What is the exact volume of a sphere with a diameter of 30 in.?

OpenStudy (gabbyalicorn):

3.14 x 30?

OpenStudy (triciaal):

volume of a sphere = 4/3 *Pi*r^3 r is the radius = 1/2 the diameter

OpenStudy (triciaal):